OpenStudy (anonymous):

approximately 2% of the nation's children (about 1.7 million) have a parent who is in jail or prison. Let X be the number of children that have an incarcerated parent from a random sample of 100 children. a) verify that this is a binomial setting and write it in notation. b) Describe what P(X=0) means in content. c) Find P(X=0) and P(X=1). d) What is the probability that two or more of the children have a parent in jail or prison?

OpenStudy (kropot72):

Do you know the conditions required for a binomial distribution?

OpenStudy (anonymous):

It has to equal to 1

OpenStudy (anonymous):

a) P=0.02 N=100

OpenStudy (kropot72):

(a) The following conditions must be met for a binomial distribution to apply: 1.The number in the sample must be fixed. n = 100 in this case. 2. Each trial must have only two possible outcomes. In this case either a parent is incarcerated or is not. 3. The probability of success on any one trial equals a value p which is constant throughout the trials. In this case p = 0.02. 4. The trial are independent of one another. Al the above conditions are met, so we can write: X ~ Bin(100, 0.02)

OpenStudy (kropot72):

(b) Do you know what P(X=0) means in the content of the question?

OpenStudy (kropot72):

P(X=0) means the probability that none of the children in the sample of 100 has an incarcerated parent. (c) \[P(X=0)=(1-0.02)^{100}=you\ can\ calculate\] \[P(X=1)=\left(\begin{matrix}100 \\ 1\end{matrix}\right)0.02^{1} \times0.98^{99}=100\times0.02\times0.98^{99}=you\ can\ calculate\]

OpenStudy (anonymous):

ThaThanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

can you help with two more?

OpenStudy (kropot72):

Sorry, I have to log out now. Please post as new questions.